Science Report – Nov 10th

Science Report 10 November 2019
SOL: 0
Crew 215 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler

Science Dome Operations: The Science Dome has been activated and is awaiting EVAs for sample acquisition and testing.

Green Hab: The green hab is fully operational and sown plants, seedlings and established plants are being maintained. See GreenHab report.

Astronomy Dome Operations: No astronomical observations have been planned for this crew rotation.

RAM Operations: The RAM Dome has been activated and is ready for contingencies.

EVA: The first EVA request for geological reconnaissance has been sent. See EVA Request.

EVA Suit Maintenance: The crew engineer will troubleshoot Suit 10 to enable repairs. Suit 2 is experiencing a reduction in airflow and will also be evaluated. See Operations Report.

Additional activities: Nil.

Copyright © The Mars Society. All rights reserved. | Main Site